# RS2-SAM2: Customized SAM2 for Referring Remote Sensing Image Segmentation

### v3

@@ Abstract
Referring Remote Sensing Image Segmentation (RRSIS) aims to segment target objects in remote sensing (RS) images based on textual descriptions. Although Segment Anything Model 2 (SAM2) has shown remarkable performance in various segmentation tasks, its application to RRSIS presents several challenges, including understanding the text-described RS scenes and generating effective prompts from text descriptions. To address these issues, we propose RS2-SAM2, a novel framework that adapts SAM2 to RRSIS by aligning the adapted RS features and textual features, providing pseudo-mask-based dense prompts, and enforcing boundary constraints. Specifically, we employ a union encoder to jointly encode the visual and textual inputs, generating aligned visual and text embeddings as well as multimodal class tokens. A bidirectional hierarchical fusion module is introduced to adapt SAM2 to RS scenes and align adapted visual features with the visually enhanced text embeddings, improving the model’s interpretation of text-described RS scenes. To provide precise target cues for SAM2, we design a mask prompt generator, which takes the visual embeddings and class tokens as input and produces a pseudo-mask as the dense prompt of SAM2. Experimental results on several RRSIS benchmarks demonstrate that RS2-SAM2 achieves state-of-the-art performance.
@@ Introduction
First, while SAM2 demonstrates strong performance on natural images, its effectiveness diminishes in remote sensing scenes due to the low distinguishability of targets and the limited contrast between foreground and background regions. Furthermore, SAM2 struggles to effectively utilize spatial information derived from textual descriptions, as illustrated in Fig. 1 (a). Although existing studies have attempted to address these issues, significant room for improvement remains. For example, SAM2-Adapter (Chen et al. 2024) modifies SAM2’s encoding process to better adapt to complex scenes. However, this unimodal approach lacks hierarchical information interaction, resulting in an insufficient fine-grained understanding of textual information. Consequently, achieving effective alignment of vision-language information and improving generalization to complex remote sensing scenes remain critical challenges in adapting SAM2 for RRSIS. Second, SAM2’s inherent lack of textual prompt integration limits its ability to generate prompts that accurately align with given textual descriptions. Previous RRSIS methods, such as RMSIN (Liu et al. 2024b), rely on independent encoding for visual and textual inputs, but such traditional encoder-decoder structures fail to effectively integrate textual information into SAM2. EVF-SAM (Zhang et al. 2024) addresses this by jointly encoding vision and language features and generating sparse prompts through a multilayer perceptron (MLP), achieving promising results on natural images. However, in remote sensing scenes, sparse prompts alone are insufficient for fine-grained local image understanding, particularly for recognizing subtle and less distinguishable objects. Thus, designing effective prompts based on text description to guide the decoding process remains a critical challenge for adapting SAM2 to RRSIS.
@@
Referring Remote Sensing Image Segmentation. RRSIS emerges as an important research direction in remote sensing, allowing users to extract geospatial information through natural language queries. Compared to the remote sensing visual grounding (RSVG) (Sun et al. 2022; Zhan, Xiong, and Yuan 2023; Lan et al. 2024) task, RRSIS focuses on fine-grained pixel-level analysis rather than region-level identification. However, research in this area remains in its early stages with limited exploration. Yuan et al. (Yuan et al. 2024) pioneer this task by introducing the RefsegRS dataset and adapting the LAVT (Yang et al. 2022) framework for remote sensing. To address challenges like small and fragmented targets easily confused with complex backgrounds, they develop a multi-level feature fusion mechanism incorporating linguistic guidance, significantly enhancing small object detection. Subsequently, Liu et al. (Liu et al. 2024b) build upon the RSVG dataset DIOR-RSVG (Zhan, Xiong, and Yuan 2023) to establish the first large-scale RRSIS dataset, RRSIS-D. They further introduce the RMSIN architecture, which leverages rotational convolution operations to better handle spatial-scale variations and directional complexities in remote sensing images, providing a more robust modeling framework for RRSIS. Building upon RMSIN (Liu et al. 2024b), Lei et al. propose FIANet (Lei et al. 2024), which focuses more on fine-grained vision-language interaction and adaptive understanding of objects at different scales.
@@
Segment Anything Model. SAM (Kirillov et al. 2023) is an interactive model designed to generate non-semantic segmentation masks based on a variety of input prompts. Leveraging large-scale training data, it demonstrates strong generalization across diverse segmentation tasks. SAM has also been widely utilized in multiple domains, including remote sensing image analysis (Wang et al. 2024), video object tracking (Cheng et al. 2023; Yan et al. 2024), and medical image segmentation (Yue et al. 2024; Cheng et al. 2024). To improve its accuracy and efficiency, several optimized versions (Zhang et al. 2023; Xiong et al. 2024; Zhong et al. 2024) have been developed. More recently, SAM2 (Ravi et al. 2024) has been introduced for video segmentation (Rong et al. 2025; Lan et al. 2023), incorporating multi-scale feature encoding to enhance segmentation robustness. Despite these advances, SAM remains limited by its lack of linguistic understanding, preventing it from directly handling referring segmentation tasks that require textual guidance. To address this limitation, recent works (Lai et al. 2024; Xu et al. 2023; Rasheed et al. 2024) have explored the integration of multimodal large language models (MLLMs) to enhance SAM’s ability to process language-based instructions. LISA (Lai et al. 2024), for example, fine-tunes LLaVA (Liu et al. 2024a) to derive hidden embeddings, thereby producing multimodal features for enhanced segmentation. u-LLaVA (Xu et al. 2023) extends this paradigm by enabling simultaneous region- and pixel-level segmentation through multi-task learning. Due to its lightweight pre-fusion strategy, EVF-SAM (Zhang et al. 2024) leverages joint visual-language encoding to enhance text-driven segmentation prompts, thereby achieving superior segmentation accuracy. However, while these models have demonstrated strong performance on natural images, their effectiveness in complex remote sensing scenes remains limited, highlighting the need for further refinement and adaptation.
@@
The architecture of the proposed RS2-SAM2 framework is depicted in Fig. 2. RS2-SAM2 comprises four essential parts: the union encoder, the bidirectional hierarchical fusion module, the mask prompt generator, and the SAM2 model. Given an input remote sensing image $\mathcal{I}$ and its associated textual description =$\mathcal{E}=\left\{e_{l}\right\}_{l=1}^{L}$, where $L$ is the number of words, the union encoder processes both modalities to extract a multimodal [CLS] token, visual patch embeddings, and textual embeddings. The SAM2 image encoder, incorporating the bidirectional hierarchical fusion module, further refines the extracted image features by leveraging textual information. Next, the mask prompt generator utilizes the visual patch embeddings and the multimodal [CLS] token to produce a pseudo-mask estimate for the target object. This prior mask, along with the multimodal [CLS] token, serves as a guiding signal for SAM2. Ultimately, the SAM2 decoder synthesizes the extracted image features and generated prompts to produce high-precision segmentation masks.
@@
Model Settings. We initialize the key modules of SAM2 and the union encoder using pre-trained weights from SAM2-Hiera-Large (Ravi et al. 2024) and BEiT-3-Large (Wang et al. 2023). For feature extraction, each image is resized to $\times 1024$ 1024 and $\times 224$ 224, which are then fed into the SAM2 image encoder with an output dimension of $C=256$ and the union encoder with an output dimension of $D=1024$. Unlike SAM2’s full capabilities, we do not employ its memory mechanism, focusing exclusively on image processing. In the bidirectional hierarchical fusion module, we set the text weight coefficient $\alpha_{t}$ to 0.2 and the image weight coefficient $\alpha_{i}$ to 0.5.

### v4

@@ Abstract
Referring Remote Sensing Image Segmentation (RRSIS) aims to segment target objects in remote sensing (RS) images based on textual descriptions. Although Segment Anything Model 2 (SAM2) has shown remarkable performance in various segmentation tasks, its application to RRSIS presents several challenges, including understanding the text-described RS scenes and generating effective prompts from text. To address these issues, we propose RS2-SAM2, a novel framework that adapts SAM2 to RRSIS by aligning the adapted RS features and textual features while providing pseudo-mask-based dense prompts. Specifically, we employ a union encoder to jointly encode the visual and textual inputs, generating aligned visual and text embeddings as well as multimodal class tokens. A bidirectional hierarchical fusion module is introduced to adapt SAM2 to RS scenes and align adapted visual features with the visually enhanced text embeddings, improving the model’s interpretation of text-described RS scenes. To provide precise target cues for SAM2, we design a mask prompt generator, which takes the visual embeddings and class tokens as input and produces a pseudo-mask as the dense prompt of SAM2. Experimental results on several RRSIS benchmarks demonstrate that RS2-SAM2 achieves state-of-the-art performance.
Codehttps://github.com/rongfu-dsb/RS2-SAM2
@@ Introduction
First, while SAM2 performs well on natural images, its effectiveness declines in remote sensing scenes due to the low target distinguishability and limited foreground–background contrast. Moreover, SAM2 struggles to exploit spatial information from textual descriptions, as shown in Fig. 1 (a). Although existing studies have attempted to address these issues, significant room for improvement remains. For example, SAM2-Adapter (Chen et al. 2024) modifies SAM2’s encoding process to better adapt to complex scenes. However, this unimodal approach lacks hierarchical information interaction, resulting in an insufficient fine-grained understanding of textual information. Consequently, achieving effective alignment of vision-language information and improving generalization to complex remote sensing scenes remain critical challenges in adapting SAM2 for RRSIS. Second, SAM2’s lack of textual prompt integration limits its ability to generate prompts aligned with textual descriptions. Previous RRSIS methods, such as RMSIN (Liu et al. 2024b), rely on independent encoding for visual and textual inputs, but such traditional encoder-decoder structures fail to effectively integrate textual information into SAM2. EVF-SAM (Zhang et al. 2024) mitigates this by jointly encoding both modalities and generating sparse prompts via an MLP, achieving promising results on natural images. However, in remote sensing, sparse prompts alone cannot ensure fine-grained understanding, especially for subtle or indistinct objects. Designing effective text-based prompts to guide decoding therefore remains a key challenge for adapting SAM2 to RRSIS.
@@
Referring Remote Sensing Image Segmentation. RRSIS emerges as an important research direction in remote sensing (Wang et al. 2025a, b), allowing users to extract geospatial information through natural language queries (Yingrui Ji 2025). Compared to the remote sensing visual grounding (RSVG) (Sun et al. 2022; Zhan, Xiong, and Yuan 2023; Lan et al. 2024) task, RRSIS focuses on fine-grained pixel-level analysis rather than region-level identification. However, research in this area remains in its early stages with limited exploration. Yuan et al. (Yuan et al. 2024) pioneer this task by introducing the RefsegRS dataset and adapting the LAVT (Yang et al. 2022) framework for remote sensing. To address challenges like small and fragmented targets easily confused with complex backgrounds, they develop a multi-level feature fusion mechanism incorporating linguistic guidance, significantly enhancing small object detection. Subsequently, Liu et al. (Liu et al. 2024b) build upon the RSVG dataset DIOR-RSVG (Zhan, Xiong, and Yuan 2023) to establish the first large-scale RRSIS dataset, RRSIS-D. They further introduce the RMSIN architecture, which leverages rotational convolution operations to better handle spatial-scale variations and directional complexities in remote sensing images, providing a more robust modeling framework for RRSIS. Building upon RMSIN (Liu et al. 2024b), Lei et al. propose FIANet (Lei et al. 2024), which focuses more on fine-grained vision-language interaction and adaptive understanding of objects at different scales.
@@
Segment Anything Model. SAM (Kirillov et al. 2023) is an interactive model designed to generate non-semantic segmentation masks based on a variety of input prompts. Leveraging large-scale training data, it demonstrates strong generalization across diverse segmentation tasks. SAM has also been widely utilized in multiple domains, including remote sensing image analysis (Wang et al. 2024b, 2025c, a), video object tracking (Cheng et al. 2023; Yan et al. 2024), and medical image segmentation (Yue et al. 2024; Cheng et al. 2024). To improve its accuracy and efficiency, several optimized versions (Zhang et al. 2023; Xiong et al. 2024; Zhong et al. 2024) have been developed. More recently, SAM2 (Ravi et al. 2024) has been introduced for video segmentation (Rong et al. 2025; Lan et al. 2023), incorporating multi-scale feature encoding to enhance segmentation robustness. Despite these advances, SAM remains limited by its lack of linguistic understanding, preventing it from directly handling referring segmentation tasks that require textual guidance. To address this limitation, recent works (Lai et al. 2024; Xu et al. 2023; Rasheed et al. 2024) have explored integrating multimodal large language models (MLLMs) to enhance SAM’s ability to process language-based instructions. LISA (Lai et al. 2024), for example, fine-tunes LLaVA (Liu et al. 2024a) to derive hidden embeddings, thus producing multimodal features for enhanced segmentation. u-LLaVA (Xu et al. 2023) extends this paradigm by enabling simultaneous region- and pixel-level segmentation through multi-task learning. With its lightweight pre-fusion strategy, EVF-SAM (Zhang et al. 2024) leverages joint visual-language encoding to strengthen text-driven segmentation prompts, achieving superior segmentation accuracy. However, while these models demonstrate strong performance on natural images, their effectiveness in complex remote sensing scenes (Zhang et al. 2025) remains limited, underscoring the need for further refinement and adaptation.
@@
The architecture of the proposed RS2-SAM2 framework is depicted in Fig. 2. RS2-SAM2 comprises four essential parts: the union encoder, the bidirectional hierarchical fusion module, the mask prompt generator, and the SAM2 model. Given an input remote sensing image $\mathcal{I}$ and its associated textual description $\mathcal{E}=\left\{e_{l}\right\}_{l=1}^{L}$, where $L$ is the number of words, the union encoder processes both modalities to extract a multimodal [CLS] token, visual patch embeddings, and textual embeddings. The SAM2 image encoder, incorporating the bidirectional hierarchical fusion module, further refines the extracted image features by leveraging textual information. Next, the mask prompt generator utilizes the visual patch embeddings and the multimodal [CLS] token to produce a pseudo-mask estimate for the target object. This prior mask, along with the multimodal [CLS] token, serves as a guiding signal for SAM2. Ultimately, the SAM2 decoder synthesizes the extracted image features and generated prompts to produce high-precision segmentation masks.
@@
Model Settings. We initialize the key modules of SAM2 and the union encoder using pre-trained weights from SAM2-Hiera-Large (Ravi et al. 2024) and BEiT-3-Large (Wang et al. 2023). For feature extraction, each image is resized to $1024\times 1024$ and $224\times 224$, which are then fed into the SAM2 image encoder with an output dimension of $C=256$ and the union encoder with an output dimension of $D=1024$. Unlike SAM2’s full capabilities, we do not employ its memory mechanism, focusing exclusively on image processing. In the bidirectional hierarchical fusion module, we set the text weight coefficient $\alpha_{t}$ to 0.2 and the image weight coefficient $\alpha_{i}$ to 0.5.
@@ Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant 62431020, the National Key Research and Development Program of China under Grant 2024YFE0111800, and the Fundamental Research Funds for the Central Universities under Grant 2042025kf0030.